# Bibliography.bib

### v1

@@ 2. Unequally warped $\mu$-bubbles and diameter bounds
00 0 $\leqslant\int_{\partial\Omega}u^{-\gamma}\Big{[}-\frac{X^{2}}{n-1}+\frac{2\gamma}{n-1}XY-\frac{\gamma^{2}}{n-1}Y^{2}-(n-1)\lambda$
@@ 3. Unequally weighted isoperimetric profile and volume bounds
Therefore $(M,g)$ satisfies the main condition (1) (with $\lambda/(n-1)$ in place of $\lambda$), but it has $\pi_{1}(M)=\mathbb{Z}$.
Let $v\in C^{\infty}(M)$, and set $\eta:=v/u$, where $u$ is as above. Then $v=u\eta$, and integrating by parts
$\int_{M}\gamma_{0}|\nabla v|^{2}+\operatorname{Ric}v^{2}$ $=\int_{M}\gamma_{0}u^{2}|\nabla\eta|^{2}+2\gamma_{0}u\eta\langle\nabla u,\nabla\eta\rangle+\gamma_{0}\eta^{2}|\nabla u|^{2}+\operatorname{Ric}u^{2}\eta^{2}$
@@
4. Isoperimetry and volume growth under $\mathrm{Ric}\geqslant 0$ and $\mathrm{biRic}\geqslant n-2$
In this section, we finally prove Theorem 2. The following lemma uses $\mu$-bubbles to construct large sets with small boundary areas.
@@ 4. Isoperimetry and volume growth under $\mathrm{Ric}\geqslant 0$ and $\mathrm{biRic}\geqslant n-2$
Let $(M^{n},g)$ be a smooth complete noncompact Riemannian manifold with $3\leqslant n\leqslant 5$, and $S$ be a connected bounded set, such that $\mathrm{biRic}\geqslant n-2$ on $M\setminus S$. Then there is a universal constant $C$ with the following property: for any $20<\varepsilon<n-$ 2 there exists a connected, smooth, bounded region $\Omega\supset\supset S$, such that $d(\partial\Omega,S)\leqslant C\varepsilon^{-1/2}$, and each connected component $\Sigma$ of $\partial\Omega$ satisfies
$-\Delta_{\Sigma}+\mathrm{Ric}_{\Sigma}-(n-2-\varepsilon)\geqslant 0,$ (49)
@@
By mollifying, we find $q\in C^{\infty}(M)$ such that $|\nabla q|\leqslant 2$ everywhere, $q=0$ on $S$, and $q\geqslant 0$, $|q-d(\cdot,S)|\leqslant 1$ on $M\setminus S$. Set the constant $C(n)=\frac{1}{n-1}\min\big{\{}\frac{1}{5},1-\frac{5}{16}(n-1)\big{\}}$, and consider
$h=\sqrt{\varepsilon/C(n)}\cot\Big{(}\frac{1}{2}\sqrt{C(n)\varepsilon}\,q\Big{)}.$
@@
$|\nabla h|\leqslant C(n)h^{2}+\varepsilon.$
Set $\Omega_{-}:=S$ and $\Omega_{+}:=\big{\{}q<2\pi/\sqrt{C(n)\varepsilon}\big{\}}$. Similarly as in the beginning of Section 2, we may find a stable $\mu$-bubble $\widetilde{\Omega}$ associated to the energy
$E(\Omega):=|\partial^{*}\Omega|-\int_{M}(\chi_{\Omega}-\chi_{\Omega_{0}})h,$
where $\Omega_{0}$ is a fixed set with $\Omega_{-}\subset\subset\Omega_{0}\subset\subset\Omega_{+}$. By construction we have $\Omega_{-}\subset\subset\widetilde{\Omega}\subset\subset\Omega_{+}$, in particular, $d(\partial\widetilde{\Omega},S)\leqslant C\varepsilon^{-1/2}$.
Following the proof of [KaiIntermediate, Lemma 3.1], the stability inequality implies that $\partial\widetilde{\Omega}$ (hence each of its components) satisfies the spectral condition (49). Finally, we let $\Omega^{\prime}$ be the connected component of $\widetilde{\Omega}$ containing $S$, and let $\Omega$ be the union of $\Omega^{\prime}$ with all the bounded connected components of $M\setminus\Omega^{\prime}$. Note that $\partial\Omega$ is a sub-collection of the connected components of $\partial\widetilde{\Omega}$, hence all the conditions of the lemma are fulfilled. ∎
@@
Let $(M^{n},g)$ be a smooth complete noncompact Riemannian manifold with $3\leqslant n\leqslant 5$. Fix $o\in M$. Assume that $\mathrm{Ric}\geqslant 0$, $M$ has one end, and $\mathrm{biRic}\geqslant n-2$ outside a compact set $K$. Then there are $r_{0}:=r_{0}(o,M,K)>0$, $C_{0}:=C_{0}(o,M,K)>0$, and universal constants $C,L>0$ such that the following holds.
@@
    $-\Delta_{\Sigma}+\operatorname{Ric}_{\Sigma}-(n-2-\varepsilon)\geqslant 0.$ (50)
@@
    The latter diameter bound can be obtained following verbatim the argument in [ChodoshLiStryker, Lemma 2.2], which involves the almost splitting theorem of Cheeger-Colding [ChCo0, ChCo1], and a uniform diameter bound of $\partial G_{1}$, $\partial G_{2}$. When $n\geqslant 4$, this uniform diameter bound comes from (50) and the last part of Corollary 1. The Corollary is applied with the choice $\gamma=1$, thus it requires $61<\frac{4}{n-2}\ \Leftrightarrow\ n<$ 6, which is satisfied by our assumptions (we cannot apply Item 1 of Theorem 1 since the result there depends on the first eigenfunction of the relevant Schrödinger operator). For the case $n=3$, we may alternatively apply [KaiSurfaces, Theorem 1.4] to get the diameter bounds of the (surfaces) $\partial G_{1}$, $\partial G_{2}$.

### v2

@@ 2. Unequally warped $\mu$-bubbles and diameter bounds
[…]
@@
Remark 2 (No universal diameter bound).
Suppose $n\geqslant 4$ and $\frac{4}{n-1}<\gamma\leqslant\frac{n-1}{n-2}$, which is the range included in Theorem 1 but not in Corollary 1. Thus, we only obtain the diameter bound (2) depending on the maximum and minimum of $u$. Here let us show that, even assuming $M$ is closed, there is no universal diameter upper bound on $M$ (that depends only on $n,\gamma$).
Fix $L\gg 1$. Fix a cutoff function $\eta_{0}$ so that $\eta_{0}|_{[-\infty,0)}\equiv 0$ and $\eta_{0}|_{[1,\infty)}\equiv 1$. Let $\delta,\mu>0$ be constants to be chosen. Let $\eta$ be the cutoff function such that:
  1. (1)
    $\eta\equiv 0$ on $(-\infty,-L-\mu]\cup[-\delta,\delta]\cup[L+\mu,+\infty)$ and $\eta\equiv 1$ on $[-L,-2\delta]\cup[2\delta,L]$,
  2. (2)
    on each interval $[-L-\mu,-L]$, $[-2\delta,-\delta]$, $[\delta,2\delta]$, $[L,L+\mu]$, the function $\eta$ interpolates between 0 and 1 according to the model function $\eta_{0}$.
Set $\alpha=\frac{2\gamma}{n-1}$ as usual. Choose constants $a,b>0$ such that
$-\frac{a^{2}}{n-1}-\Big{(}\gamma-\frac{n-2}{n-1}\gamma^{2}\Big{)}\frac{b^{2}}{(\gamma-\alpha)^{2}}-(n-1)+ab=0.$ (25)
The existence of such numbers follows by calculating the discriminant:
[…]
$\Leftrightarrow\ \ \gamma>\frac{4}{n-1},$
and the second line is exactly our assumption. From the sign of the terms in (25), it follows that we can choose $a,b$ to be positive.
Now we let $u(r)$ solve the differential equation
[…]
The solution is positive, even, and smooth on the entire $\mathbb{R}$. Also, note that $u\equiv 1$ in $[-\delta,\delta]$, and $u$ is constant on $(-\infty,-L-\mu]\cup[L+\mu,+\infty)$. Then let $h=h(r)$ solve
[…]
Since the solution is an odd function, below we only discuss within the range $r\geqslant 0$ (and the case $r<0$ follows by symmetry). We claim that there is a choice of $\delta$, such that the solution of (26) (27) satisfies $h(2\delta)u(2\delta)^{\alpha-\gamma}=-a$. To see this, we set a new function $Q=hu^{\alpha-\gamma}$, and thus (27) is converted to
[…]
When $\delta\to 0$ the solution satisfies $Q(2\delta)\to 0$. On the other hand, note that $\eta=0$ on $[0,\delta]$, thus in this interval we have $Q^{\prime}=-\frac{Q^{2}}{n-1}-(n-1)$. Thus there exists $\delta_{0}>0$ such that, when $\delta\nearrow\delta_{0}$ we have $Q(2\delta)\to-\infty$. By continuity, there exists a $\delta$ as claimed.
Next, note that $Q\equiv-a$ is a solution of (28) in $[2\delta,L]$, due to our choice (25) and $\eta\equiv 1$. By uniqueness of ODE, for our specific choice of $\delta$ the solution of (28) must satisfy $Q\equiv-a$ on $[2\delta,L]$. Thus for (27), we have $h=-au^{\gamma-\alpha}$ on $[2\delta,L]$.
Next, by continuity, there is a sufficiently small $\mu$ such that the solution of (27) exists on $[-L-\mu,L+\mu]$. When $r\geqslant L+\mu$, (27) becomes
$h^{\prime}u(L+\mu)^{\alpha-\gamma}=-\frac{u(L+\mu)^{2\alpha-2\gamma}}{n-1}h^{2}-(n-1),$
thus the solution is
$h(r)=-(n-1)u(L+\mu)^{-(\alpha-\gamma)}\cot(r_{0}-r),$
for some $r_{0}>L+\mu$. To summarize, the equation (27) has a solution $h=h(r)$ on a maximal interval $(-r_{0},r_{0})\supset[-L,L]$, and $h$ blows up like $-(n-1)u(L+\mu)^{-(\alpha-\gamma)}\cot|r_{0}\mp r|$ near the endpoints ±$\pm r_{0}$.
Finally, we let $f(r)$ solve the ODE
[…]
This equation, along with (27), come from analyzing the equality case in the computations from (15) to (17). Note that $f$ is even, since $u$ is even and $h$ is odd. When $r\geqslant L+\mu$, the ODE becomes
$(n-1)\frac{f^{\prime}(r)}{f(r)}=-(n-1)\cot(r_{0}-r)\ \ \Rightarrow\ \ f(r)=c\sin(r_{0}-r)\ \ \text{for some $c>0$.}$
For $\varepsilon\ll 1$ consider the metric $g=\mathrm{d}r^{2}+\varepsilon^{2}f(r)^{2}g_{\mathbb{S}^{n-1}}$. By the asymptotic of $f$ near ±$\pm r_{0}$, this represents a metric on $\mathbb{S}^{n}$ with acute cone singularities at the two poles $r=\pm r_{0}$. Note that $\operatorname{Ric}(\partial_{r},\partial_{r})=-(n-1)f^{\prime\prime}/f$, and $\operatorname{Ric}(e,e)$ (where $e$ is tangent to $\mathbb{S}^{n-1}$) grows like $r^{-2}$ near the poles. So we may choose $\varepsilon$ so small such that $\operatorname{Ric}(\partial_{r},\partial_{r})$ is the minimal eigenvalue of $\operatorname{Ric}$ (see, e.g., the argument at the beginning of Remark 4).
From the ODEs (27), (29) it can be directly verified that
$-\gamma\Big{[}\frac{u^{\prime\prime}}{u}+(n-1)\frac{f^{\prime}}{f}\frac{u^{\prime}}{u}\Big{]}-(n-1)\frac{f^{\prime\prime}}{f}=(n-1).$
This implies $-\gamma\Delta u+u\operatorname{Ric}(\partial_{r},\partial_{r})=(n-1)u$.
Since $g$ is conic and $u$ is constant near the pole, we can smooth the metric near the pole, while not decreasing the minimal eigenvalue of $\operatorname{Ric}$. As a consequence, we still have $-\gamma\Delta u+\operatorname{Ric}u\geqslant(n-1)u$ in the smoothed metric. This implies $\lambda_{1}(-\gamma\Delta+\operatorname{Ric})\geqslant n-1$. On the other hand, note that $\operatorname{diam}(M,g)>2L$, which can be made arbitrarily large.
@@ 3. Unequally weighted isoperimetric profile and volume bounds
Therefore $(M,g)$ satisfies the main condition (1) (with $\frac{\lambda}{n-1}$ in place of $\lambda$), but it has $\pi_{1}(M)=\mathbb{Z}$. Let $v\in C^{\infty}(M)$, and set $\eta:=v/u$, where $u$ is as above. Then $v=u\eta$, and integrating by parts
$\int_{M}\gamma_{0}|\nabla v|^{2}+\operatorname{Ric}v^{2}$ $=\int_{M}\gamma_{0}u^{2}|\nabla\eta|^{2}+2\gamma_{0}u\eta\langle\nabla u,\nabla\eta\rangle+\gamma_{0}\eta^{2}|\nabla u|^{2}+\operatorname{Ric}u^{2}\eta^{2}$ (52)
@@
Remark 5 (Smoothing the eigenfunctions).
This remark is to prove what claimed in Remark 1 in the introduction. Given $\lambda>0$ and $0\leqslant\gamma\leqslant\frac{n-1}{n-2}$, assume the positivity of Dirichlet energy
$\lambda_{1}(-\gamma\Delta+\operatorname{Ric})\geqslant(n-1)\lambda,$ (54)
on the $n$-dimensional manifold $M$. We first aim at proving the following:
If $M$ is compact and (54) holds, then (55)
                                 $\operatorname{vol}(\widetilde{M})\leqslant\lambda^{-n/2}\operatorname{vol}(\mathbb{S}^{n})$, and $\pi_{1}(M)$ is finite.
By regularizing the function $\mathrm{Ric}$, for each $0<\varepsilon\ll\lambda$ there exists $V_{\varepsilon}\in C^{\infty}(M)$ such that $\|V_{\varepsilon}-\mathrm{Ric}\|_{C^{0}}\leqslant\varepsilon$. Note that $\lambda_{1}(-\gamma\Delta+V_{\varepsilon})\geqslant(n-1)\lambda-\varepsilon$. Let $u_{\varepsilon}\in C^{\infty}(M)$ be the first eigenfunction of $-\gamma\Delta+V_{\varepsilon}$. Notice that $u_{\varepsilon}$ is positive, and we have
$-\gamma\Delta u_{\varepsilon}+\operatorname{Ric}u_{\varepsilon}\geqslant\Big{[}(n-1)\lambda-2\varepsilon\Big{]}u_{\varepsilon}.$ (56)
Thus, applying Theorem 1 with (56) and taking $\varepsilon\to 0$, we have proved the desired (55).
Let us now deal with the rigidity part. For simplicity, after rescaling, assume $\lambda=1$. Hence, assume we have $\lambda_{1}(-\gamma\Delta u+\operatorname{Ric})\geqslant n-1$ and $\operatorname{vol}(M)=\operatorname{vol}(\mathbb{S}^{n})$. Let us show that $M$ is isometric to the standard sphere of radius $1$.
Let $u\in W^{1,2}(M)$ be the first eigenfunction of $-\gamma\Delta+\mathrm{Ric}$, so that $-\gamma\Delta u+\operatorname{Ric}u=\lambda u$ for some $\lambda\geqslant(n-1)$. Normalize $u$ so that $\int_{M}u^{2}=1$. Set $\lambda_{\varepsilon}:=\lambda_{1}(-\gamma\Delta+V_{\varepsilon})$, and let $u_{\varepsilon}\in C^{\infty}(M)$ be the corresponding first eigenfunction, normalized so that $\int_{M}u_{\varepsilon}^{2}=1$. Note that $|\lambda-\lambda_{\varepsilon}|\leqslant\varepsilon$, and $u,u_{\varepsilon}$ are all positive.
Let us show that $u_{\varepsilon}\to u$ in $C^{1,\beta}$ for every $10<\beta<$ 1. Note that $\int_{M}|\nabla u_{\varepsilon}|^{2}$ are uniformly bounded, so there is a subsequence converging to some $\widetilde{u}\in W^{1,2}$ strongly in $L^{2}$ and weakly in $W^{1,2}$. Using lower semi-continuity, and the fact that $u_{\varepsilon}$ are normalized eigenfunctions, we have:
$\int_{M}\gamma|\nabla\widetilde{u}|^{2}+\operatorname{Ric}\widetilde{u}^{2}\leqslant\liminf_{\varepsilon\to 0}\int_{M}\gamma|\nabla u_{\varepsilon}|^{2}+\lim_{\varepsilon\to 0}\int_{M}V_{\varepsilon}u_{\varepsilon}^{2}\leqslant\lim_{\varepsilon\to 0}\lambda_{\varepsilon}=\lambda.$
So it must happen that $\widetilde{u}=u$. Thus $u_{\varepsilon}\to u$ in $L^{2}$. Now, applying standard elliptic regularity, this can be bootstrapped to imply $u_{\varepsilon}\to u$ in $W^{2,p}$ for every $p>1$, hence in $C^{1,\beta}$ for every $10<\beta<$ 1.
By Theorem 1 (specifically, (50) in its proof) and the rigidity assumption, we have
$\operatorname{vol}(M)$ $\leqslant\min(u_{\varepsilon})^{-\alpha}\int_{M}u_{\varepsilon}^{\alpha}\leqslant\operatorname{vol}(\mathbb{S}^{n})\big{(}1+o_{\varepsilon}(1)\big{)}=\operatorname{vol}(M)\big{(}1+o_{\varepsilon}(1)\big{)}.$
Passing this to the limit, we get
$\operatorname{vol}(M)\leqslant\min(u)^{-\alpha}\int_{M}u^{\alpha}\leqslant\operatorname{vol}(M).$
Hence $u$ is constant, and we are reduced to pointwise Ricci bound, from which the rigidity follows from the classical Bishop-Gromov volume comparison.
Finally, the Bonnet-Myers result (2) follows from (56) and the $C^{1,\beta}$-convergence obtained above. ∎
4. Isoperimetry under $\mathrm{Ric}\geqslant 0$ and spectral $\mathrm{biRic}\geqslant n-2$
We aim at showing that when $3\leqslant n\leqslant 5$, a spectral biRicci condition descends to a spectral Ricci condition on properly chosen warped $\mu$-bubbles. We will keep track of the sharp constants in the process, and we will use the forthcoming computations to show Lemma 5. We stress that the following computations are inspired by [CLMS, Theorem 4.3]. Similar computations have appeared in [Mazet, Theorem 4.1] in the proof of the stable Bernstein problem in $\mathbb{R}^{6}$, while we were revising the final draft of this paper.
Let $(M^{n},g)$ be a complete smooth Riemannian manifold. Let $3\leqslant n\leqslant 5$, and $0\leqslant\gamma<6-n$. It will also be clear from the computations that when $n=4$ we can choose $0\leqslant\gamma\leqslant 2$. Let $u$ be a positive smooth function on $M$ such that, for some $\lambda\in\mathbb{R}$,
$-\gamma\Delta u+u\mathrm{biRic}\geqslant\lambda u.$ (57)
For an arbitrary fixed domain $\Omega_{0}$ with $\Omega_{-}\subset\subset\Omega_{0}\subset\subset\Omega_{+}\subset M$, consider the functional
$E(\Omega):=\int_{\partial^{*}\Omega}u^{\gamma}-\int(\chi_{\Omega}-\chi_{\Omega_{0}})hu^{\gamma},$ (58)
defined on sets of finite perimeter, where $\partial^{*}\Omega$ is the reduced boundary of $\Omega$. Take $\Omega$ to be a minimizer, and notice $\partial\Omega$ is smooth because $3\leqslant n\leqslant 5$.
$0=\frac{\mathrm{d}}{\mathrm{d}t}E(\Omega_{t})\Big{|}_{t=0}=\int_{\partial\Omega}\big{(}H+\gamma u^{-1}u_{\nu}-h\big{)}u^{\gamma}\varphi.$ (59)
Since $\varphi$ is arbitrary we have $H=h-\gamma u^{-1}u_{\nu}$. Then, computing the second variation,
00 0 $\leqslant\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}E(\Omega_{t})\Big{|}_{t=0}$
$=\int_{\partial\Omega}\Big{[}-\Delta_{\partial\Omega}\varphi-|\text{II}|^{2}\varphi-\operatorname{Ric}(\nu,\nu)\varphi-\gamma u^{-2}u_{\nu}^{2}\varphi$
$\hskip 72.0pt+\gamma u^{-1}\varphi\big{(}\Delta u-\Delta_{\partial\Omega}u-Hu_{\nu}\big{)}-\gamma u^{-1}\langle\nabla_{\partial\Omega}u,\nabla_{\partial\Omega}\varphi\rangle-h_{\nu}\varphi\Big{]}u^{\gamma}\varphi.$
Set $\varphi:=\psi u^{-\gamma/2}$, so $u^{\gamma}\varphi^{2}=\psi^{2}$. Then we get
00 0 $\leqslant\int_{\partial\Omega}\Big{[}-\Delta_{\partial\Omega}(\psi u^{-\gamma/2})\psi u^{\gamma/2}-\gamma u^{-1}\psi^{2}\Delta_{\partial\Omega}u-\gamma u^{\gamma/2-1}\psi\langle\nabla_{\partial\Omega}u,\nabla_{\partial\Omega}(\psi u^{-\gamma/2})\rangle\Big{]}$ (60)
$\qquad\qquad+\psi^{2}\Big{[}-|\text{II}|^{2}-\operatorname{Ric}(\nu,\nu)-\gamma u^{-2}u_{\nu}^{2}+\gamma u^{-1}\big{(}\Delta u-Hu_{\nu}\big{)}-h_{\nu}\Big{]}$
$=:P+Q.$
Recall that, see [CLMS, Page 13],
$|\text{II}|^{2}+\mathrm{Ric}(\nu,\nu)\geqslant\mathrm{biRic}-\mathrm{Ric}_{\partial\Omega}+\frac{6-n}{4}H^{2}.$ (61)
Let us call $Y=u^{-1}u_{\nu}$, thus $H=h-\gamma Y$. By (61) and (57) we have:
$-|\text{II}|^{2}-\mathrm{Ric}(\nu,\nu)+\gamma u^{-1}\Delta u$ $\leqslant-\mathrm{biRic}+\mathrm{Ric}_{\partial\Omega}-\frac{6-n}{4}H^{2}+\gamma u^{-1}\Delta u$
$\leqslant-\lambda+\mathrm{Ric}_{\partial\Omega}-\frac{6-n}{4}H^{2}.$
Thus
$Q$ $\leqslant\int_{\partial\Omega}\psi^{2}\Big{[}-\lambda+\mathrm{Ric}_{\partial\Omega}-\frac{6-n}{4}(h-\gamma Y)^{2}-\gamma Y^{2}-\gamma(h-\gamma Y)Y+|\nabla h|\Big{]}=$ (62)
$\leqslant\int_{\partial\Omega}\psi^{2}\Big{[}-\lambda+\mathrm{Ric}_{\partial\Omega}+|\nabla h|-\frac{6-n}{4}h^{2}+\frac{4-n}{2}\gamma hY+\left(\frac{n-2}{4}\gamma^{2}-\gamma\right)Y^{2}\Big{]}.$
Notice that $6-n\leqslant\frac{4}{n-2}$ for all $n>2$, with equality if and only if $n=4$. Hence, since $\gamma<6-n$, we also have $\gamma<\frac{4}{n-2}$. It can computed directly that
$\frac{6-n}{4}h^{2}-\frac{4-n}{2}\gamma hY+\left(\gamma-\frac{n-2}{4}\gamma^{2}\right)Y^{2}$ (63)
$\quad=\left(\frac{6-n}{4}-\frac{(4-n)^{2}\gamma}{4\left(4-(n-2)\gamma\right)}\right)h^{2}+\left(\gamma-\frac{n-2}{4}\gamma^{2}\right)\left(\frac{(4-n)}{4-(n-2)\gamma}h-Y\right)^{2}.$
Denote
$c(n,\gamma):=\frac{6-n}{4}-\frac{(4-n)^{2}\gamma}{4(4-(n-2)\gamma)},$ (64)
and notice that $\gamma<6-n\Leftrightarrow c(n,\gamma)>0$, and $c(n,\gamma)=0$ if and only if $\gamma=6-n$. Moreover, since $0\leqslant\gamma<\frac{4}{n-2}$ as noticed above, we have $\gamma-\frac{n-2}{4}\gamma^{2}\geqslant 0$. Joining (62), and (LABEL:eqn:AlgebraicInequality) we thus conclude
$Q\leqslant\int_{\partial\Omega}\psi^{2}\left[-\lambda+\mathrm{Ric}_{\partial\Omega}+|\nabla h|-c(n,\gamma)h^{2}\right].$ (65)
Notice that in case $n=4$, and $\gamma=2$, (65) still holds with $c(4,2):=1/2$ (i.e., when $n=4$ we can work in the full interval $0\leqslant\gamma\leqslant 2$ as claimed above).
Then we deal with the gradient terms $P$. Notice that we have the following integration by parts formula for every smooth function $\eta\in C^{\infty}(\partial\Omega)$:
$\quad\int_{\partial\Omega}\langle\nabla_{\partial\Omega}(u^{-\gamma}\eta),\nabla_{\partial\Omega}\eta\rangle-\gamma u^{-\gamma-1}\eta^{2}\Delta_{\partial\Omega}u-\gamma u^{-1}\eta\langle\nabla_{\partial\Omega}u,\nabla_{\partial\Omega}(u^{-\gamma}\eta)\rangle$ (66)
$=\int_{\partial\Omega}u^{-\gamma}|\nabla_{\partial\Omega}\eta|^{2}-\gamma u^{-\gamma-2}\eta^{2}|\nabla_{\partial\Omega}u|^{2}.$
Thus applying (66) with $\eta=\psi u^{\gamma/2}$ we get, using Young’s inequality,
$P$ $=\int_{\partial\Omega}u^{-\gamma}|\nabla_{\partial\Omega}(\psi u^{\gamma/2})|^{2}-\gamma u^{-2}\psi^{2}|\nabla_{\partial\Omega}u|^{2}$
$=\int_{\partial\Omega}|\nabla_{\partial\Omega}\psi|^{2}+\gamma u^{-1}\psi\langle\nabla_{\partial\Omega}u,\nabla_{\partial\Omega}\psi\rangle+\left(\frac{\gamma^{2}}{4}-\gamma\right)u^{-2}\psi^{2}|\nabla_{\partial\Omega}u|^{2}$
$\leqslant\int\left(1+\frac{\gamma^{2}}{4c}\right)|\nabla_{\partial\Omega}\psi|^{2}+\left(\frac{\gamma^{2}}{4}-\gamma+c\right)u^{-2}\psi^{2}|\nabla_{\partial\Omega}u|^{2}.$
Choose $c=\gamma-\gamma^{2}/4$, which is positive since $0\leqslant\gamma\leqslant 4$ in the range we are working in. Hence we get
$S\leqslant\int\frac{4}{4-\gamma}|\nabla_{\partial\Omega}\psi|^{2}.$ (67)
Putting together (65), (67), and (60) we get
$\int_{\partial\Omega}\frac{4}{4-\gamma}\big{|}\nabla_{\partial\Omega}\psi\big{|}^{2}+\operatorname{Ric}_{\partial\Omega}\psi^{2}\geqslant\int_{\partial\Omega}\Big{[}\lambda-|\nabla h|+c(n,\gamma)h^{2}\Big{]}\psi^{2},$ (68)
for every $\psi\in C^{\infty}(\partial\Omega)$.
The previous computations can be used to infer the following key Lemma 5. In dimension $n=4$, or when $\gamma=0$, similar results have appeared in [CLMS, Theorem 4.3], and [KaiIntermediate, Lemma 3.1], respectively.
@@ 4. Isoperimetry under $\mathrm{Ric}\geqslant 0$ and spectral $\mathrm{biRic}\geqslant n-2$
Let $(M^{n},g)$ be a smooth complete noncompact Riemannian manifold with $3\leqslant n\leqslant 5$, let $0\leqslant\gamma<6-n$ (or $0\leqslant\gamma\leqslant 2$ when $n=4$), and $S$ be a compact, and connected set such that
$-\gamma\Delta+\mathrm{biRic}\geqslant n-2$ (69)
on $M\setminus S$. Then there is a constant $C:=C(n,\gamma)$ with the following property: for any $20<\varepsilon<n-$ 2 there exists a connected, smooth, bounded region $\Omega\supset\supset S$, such that $d(\partial\Omega,S)\leqslant C\varepsilon^{-1/2}$, each connected component $\Sigma$ of $\partial\Omega$ satisfies
$-\frac{4}{4-\gamma}\Delta_{\Sigma}+\mathrm{Ric}_{\Sigma}-(n-2-\varepsilon)\geqslant 0,$ (70)
@@
Fix $20<\varepsilon<n-$ 2. From the assumption (69), and using [FischerColbrieSchoenCPAM, Theorem 1], we can find a positive $u\in C^{\infty}(M\setminus S)$ such that
$-\gamma\Delta u+u\operatorname{biRic}\geqslant(n-2-\varepsilon/2)u.$
By mollifying, we find $q\in C^{\infty}(M)$ such that $|\nabla q|\leqslant 2$ everywhere, $q=0$ on $S$, and $q\geqslant 0$, $|q-d(\cdot,S)|\leqslant 1$ on $M\setminus S$. Set the constant $C:=c(n,\gamma)$ as in (64), and consider
$h=\sqrt{\varepsilon/(2C)}\cot\Big{(}\frac{1}{2}\sqrt{(C\varepsilon)/2}\,q\Big{)}.$
@@
$|\nabla h|\leqslant Ch^{2}+\varepsilon/2.$ (71)
Set $\Omega_{-}:=S$ and $\Omega_{+}:=\big{\{}q<2\pi/\sqrt{(C\varepsilon)/2}\big{\}}$. Let us consider a warped stable $\mu$-bubble $\widetilde{\Omega}$ associated to the energy
$E(\Omega):=\int_{\partial^{*}\Omega}u^{\gamma}-\int(\chi_{\Omega}-\chi_{\Omega_{0}})hu^{\gamma},$
where $\Omega_{0}$ is a fixed set with $\Omega_{-}\subset\subset\Omega_{0}\subset\subset\Omega_{+}$. By construction we have $\Omega_{-}\subset\subset\widetilde{\Omega}\subset\subset\Omega_{+}$, in particular, $d(\partial\widetilde{\Omega},S)\leqslant C\varepsilon^{-1/2}$, with a possibly different $C$ only depending on $n,\gamma$.
Now let us exploit the computations before the present Lemma. We have (57) with $\lambda:=n-2-\varepsilon/2$ on $M\setminus S$. Notice that $\widetilde{\Omega}\subset\subset M\setminus S$, so the argument before the previous Lemma, and in particular (68) together with (71), will let us conclude that $\partial\widetilde{\Omega}$ (hence each of its components) satisfies the spectral condition (70).
Finally, we let $\Omega^{\prime}$ be the connected component of $\widetilde{\Omega}$ containing $S$, and let $\Omega$ be the union of $\Omega^{\prime}$ with all the bounded connected components of $M\setminus\Omega^{\prime}$. Note that $\partial\Omega$ is a sub-collection of the connected components of $\partial\widetilde{\Omega}$, hence all the conditions of the lemma are fulfilled. ∎
@@
Let $(M^{n},g)$ be a smooth complete noncompact Riemannian manifold with $3\leqslant n\leqslant 5$, and let $0\leqslant\gamma<6-n$ (or $0\leqslant\gamma\leqslant 2$ when $n=4$). Fix $o\in M$. Assume that $\mathrm{Ric}\geqslant 0$, $M$ has one end, and $-\gamma\Delta+\mathrm{biRic}\geqslant n-2$ outside a compact set $K$. Then there are $r_{0}:=r_{0}(o,M,K)>0$, $C_{0}:=C_{0}(o,M,K)>0$, and universal constants $C,L>0$ such that the following holds.
@@
    $-\frac{4}{4-\gamma}\Delta_{\Sigma}+\operatorname{Ric}_{\Sigma}-(n-2-\varepsilon)\geqslant 0.$ (72)
@@
    The latter diameter bound can be obtained following verbatim the argument in [ChodoshLiStryker, Lemma 2.2], which involves the almost splitting theorem of Cheeger-Colding [ChCo0, ChCo1], and a uniform diameter bound of $\partial G_{1}$, $\partial G_{2}$. When $n\geqslant 4$, this uniform diameter bound comes from (72) and the last part of Corollary 1. The Corollary is applied with the $\gamma$ there being equal to $\frac{4}{4-\gamma}$ here. This requires $\frac{4}{4-\gamma}<\frac{4}{n-2}$, which is satisfied by our assumption $\gamma<6-n$. Notice that when $n=4$, it suffices to have $\leqslant$ instead of $<$ in the previous inequalities. For the case $n=3$, we may alternatively apply [KaiSurfaces, Theorem 1.4] to get the diameter bounds of the (surfaces) $\partial G_{1}$, $\partial G_{2}$. When $n=3$, $\gamma<3$, and then $\frac{4}{4-\gamma}<4$, which is what we need to apply [KaiSurfaces, Theorem 1.4].